# Protecting Bystander Privacy via Selective Hearing in LALMs

### v1

@@ 6.2 Ablation Studies
We perform ablation studies on two design factors of SH-Bench: (i). the incorporation of the main speaker description, and (ii). the model refusal behaviour with and without “I don’t know" option.
@@ 6.2.2 Refusal Behaviour
Models after BPFT learns to refuse for open-ended questions. It is also crucial to investigate whether the model after BPFT can refuse to answer under general scenario. To this end, we convert the MCQs into open-ended questions by removing the choices, and prompt the model under selective mode to observe its behaviour. The results are summarised in Table 5. As expected, both Qwen-2.5-Omni and Step-Audio-2-mini learn to refuse when the question is about the bystander, with refusal rate of 92% and 96% respectively. The refusal rate for both models before training with BPFT was merely 45% and 48% respectively.
@@ 7 Conclusion
We propose selective hearing benchmark (SH-Bench), the first benchmark evaluating selective hearing abilities in LALMs to protect bystander privacy. Together with SH-Bench, an evaluation framework is proposed to evaluate both comprehension and bystander privacy protection, with a unified criterion, selective efficacy (SE), being proposed. Moreover, bystander privacy fine-tuning (BPFT) pipeline was proposed. Experimental results demonstrates the lack of bystander privacy protection in existing LALMs, and the effectiveness of BPFT which achieves an SE of 91.7% compared to Gemini 2.5 Pro with an SE of 75.8%.
@@ 8 Limitations
Since this is the first effort to systematically evaluate bystander privacy protection in LALMs, we only focus on single main speaker scenarios which is the most common scenarios for personal AI assistants. We believe there are more challenging scenarios such as group discussions (e.g. with AI in an open space) that can be explored in the future. Moreover, since we examined a range of omni models that can receive audio and visual inputs simultaneously, the bystander privacy could also be extended to audio-visual scenarios. As BPFT is targeted for the SH-Bench setup, we did not include multiple main speakers in the training pipeline and hence the model may fail when there are more than one main speaker.
@@ Acknowledgments
We thank Dongcheng Jiang, Xiaodong Wu, and Yifan Xu for their valuable insights on methods for collecting real scenario audios. This project is funded by Trinity College, Cambridge.
@@ A.2 Accuracy on Real and Synthetic Splits
The accuracies on real and synthetic scenarios are in Table. LABEL:tab:real_vs_synth. Under both modes, the difference between real and synthetic scenario is negligible for main speakers. The main different is observed on bystanders. The bystander accuracy in the real scenario is always lower than that in the synthetic scenario under general mode since real scenarios are more challenging. Under the selective mode, models before BPFT are more likely to choose IDK on real scenario than on synthetic ones. The accuracies of models after BPFT on synthetic scenario are close to 100% and are clearly higher than real scenarios, as the training enabled models to distinguish the bystander in synthetic scenarios more easily.

### v2

@@ 6.2 Ablation Studies
We perform ablation studies on two design factors of SH-Bench: i) the incorporation of the main speaker description; and ii) the model refusal behaviour with and without “I don’t know" option.
@@ 6.2.2 Refusal Behaviour
Models after BPFT learns to refuse for open-ended questions. It is also crucial to investigate whether the model after BPFT can refuse to answer under general scenario. To this end, we convert the MCQs into open-ended questions by removing the choices, and prompt the model under selective mode to observe its behaviour. The results are summarised in Table 5. As expected, both Qwen-2.5-Omni and Step-Audio-2-mini learn to refuse when the question is about the bystander, with refusal rates of 92% and 96% respectively. The refusal rate for both models before training with BPFT was merely 45% and 48% respectively.
@@ 7 Conclusion
We propose selective hearing benchmark (SH-Bench), the first benchmark evaluating selective hearing abilities in LALMs to protect bystander privacy. Together with SH-Bench, an evaluation framework is proposed to evaluate both comprehension and bystander privacy protection, with a unified criterion, selective efficacy (SE), being proposed. Moreover, bystander privacy fine-tuning (BPFT) pipeline was proposed. Experimental results demonstrates the lack of bystander privacy protection in existing audio LLMs, and the effectiveness of BPFT which achieves an SE of 91.7% compared to Gemini 2.5 Pro with an SE of 75.8%.
@@ 8 Limitations
Since this is the first effort to systematically evaluate bystander privacy protection in audio LLMs, we only focus on single main speaker scenarios which is the most common scenarios for personal AI assistants. We believe there are more challenging scenarios such as group discussions (e.g. with AI in an open space) that can be explored in the future. Moreover, since we examined a range of omni models that can receive audio and visual inputs simultaneously, the bystander privacy could also be extended to audio-visual scenarios. As BPFT is targeted for the SH-Bench setup, we did not include multiple main speakers in the training pipeline and hence the model may fail when there are more than one main speaker.
@@ Acknowledgments
We thank Dongcheng Jiang, Xiaodong Wu, and Yifan Xu for their valuable insights on methods for collecting real scenario audios. This project is funded by Trinity College, Cambridge; and by the Generalitat Valenciana’s project PROMETEO CIPROM/2023/23.
@@ Appendix A Appendix
Models Main Speaker %Acc (Real/Synth) Bystander %Acc (Real/Synth)
General Selective General Selective
Gemini 2.5 Pro 97.6 / 97.0 97.4 / 96.6 54.4 / 76.6 68.2 / 50.2
Qwen-2.5-Omni 7B 95.6 / 96.4 95.6 / 95.4 41.4 / 55.0 53.0 / 42.2
Step-Audio-2-mini 94.8 / 93.6 94.0 / 93.4 51.0 / 58.6 35.8 / 27.2
Qwen-2.5-Omni 7B + BPFT 91.6 / 95.0 90.8 / 94.6 74.4 / 89.6 90.0 / 99.6
Step-Audio-2-mini + BPFT 97.2 / 97.6 91.6 / 97.0 73.8 / 88.2 92.8 / 99.4
Table 6: Detailed accuracies on real and synthetic scenarios.
@@ A.2 Accuracy on Real and Synthetic Splits
The accuracies on real and synthetic scenarios are in Table 6. Under both modes, the difference between real and synthetic scenario is negligible for main speakers. The main different is observed on bystanders. The bystander accuracy in the real scenario is always lower than that in the synthetic scenario under general mode since real scenarios are more challenging. Under the selective mode, models before BPFT are more likely to choose IDK on real scenario than on synthetic ones. The accuracies of models after BPFT on synthetic scenario are close to 100% and are clearly higher than real scenarios, as the training enabled models to distinguish the bystander in synthetic scenarios more easily.